OpenStudy (rjtrain1):

Which of the following number sentences illustrates the associative property of multiplication? (3 × 8) × 6 = 3 × (8 × 6) 2 × (1 × 9) = (2 × 1) × (2 × 9) 1 × 15 = 15 8 × 9 = 9 × 8 will medal and fan

NvidiaIntely (nvidiaintely):

option A.

NvidiaIntely (nvidiaintely):

yw

NvidiaIntely (nvidiaintely):

BTW @rjtrain1, you should understand this. Check out this link: https://www.google.com/search?q=associative&oq=associative&aqs=chrome..69i57j0l5.434j0j1&sourceid=chrome&ie=UTF-8

OpenStudy (rjtrain1):

btw I did not understand that link

OpenStudy (rjtrain1):

thank you anyway

OpenStudy (andrewishere):

The associative property basically means that no matter what order your variable are arranged to be multiplied, they will give you the same answer. So say 3 = a 8 = b and 6 = c (a*b)*c equals abc (a*c)*b = acb, or abc (c*b)*a = cba, or abc

NvidiaIntely (nvidiaintely):

Thanks @AndrewIsHere. Good explanation.

OpenStudy (andrewishere):

Same as with 3, 8 and 6 (3*8)*6 = 144 as 3*(8*6) also equal 144